The Three Sisters Lumber Company is considering buying a machine that planes lumber to the correct thickness. The machine is advertised to produce “6-inch lumber” having a thickness that is normally distributed, with a mean of 6 inches and a standard deviation of 0.1 inch. a. If building standards in the industry require a 99% chance of a board being between 5.85 and 6.15

1
Expert's answer
2021-07-16T08:20:37-0400

Mean = 6 inch

Standard Deviation = 0.1 inch

Standard error = "\dfrac{sd}{\sqrt n}=\dfrac{0.1}{\sqrt 1}=0.1"

"\alpha = 1-0.99=0.01"

Critical Value "= \pm 2.575"


Margin of Error = Critical Value "\times" Standard Error = "\pm 2.575\times 0.1=\pm 0.2575"


99% Confidence interval = Sample mean "\pm" Margin of error

= 6 "\pm" 0.2575 = (5.7425 , 6.2575)

Hence,

No, The Three Sisters Lumber Company should not purchase this machine because the 99% confidence interval goes beyond the industry standard.